OpenStudy (anonymous):

write equation in function notation. 6xsquared+2y=2

OpenStudy (anonymous):

move x term to the right, so 2y = 2-6x^2. Then divide by 2 on both sides, so y = f(x) = 1-3x^2. hope this helps.

OpenStudy (anonymous):

how bout 3x^2-1